# Quantum thermocouples: nonlocal conversion and control of heat in semiconductor nanostructures

### v1

@@ V Heat currents and cooling
The central part of this review has been devoted to nanoscale configurations able to generate electrical power in an isothermal conductor using some resource from a third one (usually heat). However, the same working principles have been proposed to operate these systems for a variety of different useful tasks. These include the non-local cooling of a cold reservoir, either using thermal (absorption refrigerators) or voltage (Peltier refrigerators) biases, and thermal management operations aimed at achieving control over heat flows [267]. We dedicate this last section to overview recent developments in nonlocal absorption and Peltier refrigerators and thermal management operation exclusive of multiterminal configurations: the circulation of heat and heat transistors.
@@ V.1 Absorption refrigerators
A multiterminal nanoscale system can work as an absorption refrigeration when there are at least two different temperatures among the electron reservoirs, so that under appropriate conditions the heat injected by the hottest terminal, $J_{H}<0$, forces the extraction of heat from the coldest one, $J_{C}>0$, in the absence of any electrochemical potential bias. In a three-terminal configuration with three temperatures, $T_{H}>T_{0}>T_{C}$, energy conservation ($J_{H}+J_{0}+J_{C}=0)$ and the second law of thermodynamics ($-J_{H}/T_{H}-J_{0}/T_{0}-J_{C}/T_{C}\geq 0)$ impose that heat is dissipated in the reservoir with intermediate temperature, so we can identify $J_{H}$ as the only resource and define a coefficient of performance (COP) as: $\text{COP}=J_{C}/J_{H}$ , which is bounded from above by $\text{COP}_{C}=T_{C}/(T_{0}-T_{C})(1-T_{0}/T_{H})$.
@@
FIG. 6: Phase coherent absorption refrigerator. Heat is extracted from the cold reservoir of a three terminal system as the one depicted in Figure 3(a), with scattering regions 1 and 2 being quantum dots with resonance energies $\varepsilon_{1}$ and $\varepsilon_{2}$ and broadening $\Gamma=0.2k_{B}T$. (a) Cooling power, $J_{C}=J_{L}$, as a function of the resonance energies produced by a hot tip with $\Delta T_{H}/T=1$, and isothermal terminals $L$ and $R$, $T_{L}=T_{R}=T$. (b) Cooling power as a function of the heating at the tip, $\Delta T_{H}$, and of the right terminal, $\Delta T_{R}$, for $T_{L}=T$ and $(\epsilon_{1}-\mu)/k_{B}T=2$, $(\epsilon_{2}-\mu)/k_{B}T=4$. (c) Coefficient of performance for fixed $\Delta T_{H}=1$ but otherwise same parameters as in (b). Fixed parameters: $\mu=20k_{B}T$, tip transmission $\epsilon=0.5$, $d=2.5l_{0}$, $x=1.7l_{0}$, being $l_{0}=\hbar/\sqrt{8mk_{\text{B}}T}$, with the electron mass $m$.
@@
So far, Ref. [268] is the only absorption refrigerator in a noninteracting conductor to the best to our knowledge. The reason may be that this effect is very sensitive to the elastic coupling between terminals $L$ and $R$. We verified for this review that the same mechanism works for the fully phase-coherent implementation of the noninteracting resonant tunneling thermocouple discussed in Sec. III.2 and sketched in Fig. 3(a). The tip is used as a heat source that refrigerates one of the conductor terminals depending on its position. In Fig. 6(a) we show the heat flow $J_{L}$ when the tip temperature is increased by $\Delta T_{H}$ while maintaining $T_{L}=T_{R}=T$, as a function of the resonant level energies of the quantum dots. Cooling occurs for a very specific configuration where both levels are over the electrochemical potential, $\varepsilon_{\alpha}>\mu$, and destructive interference suppresses ${\cal T}_{LR}(E)$. In contrast to the inelastic version, no refrigeration is found in the case with $\varepsilon_{\alpha}<\mu$ (note that symmetries of the transmission probabilities exchanging $\varepsilon_{1}\leftrightarrow\varepsilon_{2}$ are only expected upon moving the tip from $x$ to $d-x$). The effect is also sensitive to an increase of the temperature of the other terminal, $\Delta T_{R}$, as shown in Fig. 6(b), as expected. The coefficient of performance is shown in Fig. 6(c) as a function of this temperature, attaining only a low fraction of the Carnot limit for this non optimized choice of parameters.
@@
A more robust effect is found due to interactions. Capacitively coupled quantum dots in the arrangement of Fig. 5(a) work as absorption refrigerators when the reservoir coupled to quantum dot $g$ is either the heat injector [87] or the refrigerated cold terminal [269, 270]. In the former case, it is sufficient that the heat conversion engine discussed in Sec. IV.1 operates in a configuration in which the generated current either extracts electrons from a reservoir ($L$ or $R$) over its electrochemical potential, or injects them below it [87]. Interestingly, the refrigeration of one conductor terminal can be combined with pumping heat on the hotter (but not hotter than the gate terminal) and additionally generating power, such that all currents in the conductor are reversed [177]. In the later case (cooling the terminal coupled to $g$), their operation relies in using the hot temperature of one of the terminals coupled to $s$ to favor the same series of fluctuations, $(n_{s},n_{g})=(0,0)\rightarrow(1,0)\rightarrow(1,1)\rightarrow(0,1)\rightarrow(0,0)$, cf. Eq. (21), that was used to transfer an energy quantum $U$ from $g$ to $s$ to power the heat engine. However, this requires a non trivial choice of the system parameters that limits its operating regime, and is restricted to low cooling powers.
@@
A photonic environment can also play the role of the hot resource for absorption refrigerators. Thermal photons coupled to a double quantum dot connecting two electronic reservoirs at different temperatures can drive processes that extract heat for the colder one [271]. The inverse scenario, cooling a photonic resonator using two electronic reservoirs at different temperatures, was also explored in hybrid (superconductor-metal) configurations connected by a quantum dot [272].
@@ V.2 Nonlocal Peltier cooling
Analogously, the refrigeration side of chiral and Aharonov-Bohm engines discussed in Secs. III.2.1 and III.2, have been discussed using the split resonances of two quantum dots and taking advantage of the directionality imposed by the magnetic field [278, 279].
@@
In metallic islands, the energy balance needs to take the phonon bath into account, which adds an additional reservoir to the problem. Using the quasiparticle spectral properties of superconductor terminals in the current-carrying conductor, the temperature of a capacitively coupled single-electron box has been predicted to be reduced from $200\,\mathrm{mK}$ to $50\,\mathrm{mK}$ for islands of size $30\times 50\times 300\,\mathrm{nm^{3}}$ [284]. One can also extract heat from the phonon bath that couples to the nanostructure [285, 158, 159].
@@ V.3.1 Heat circulators
The three terminal configuration also enables the possibility of defining circulation operations, where the heat injected from one terminal flows preferably in one direction (clock- or counterclockwise) to the next one, while transport in the opposite way is partially or totally blocked. A clear case where this happens is in chiral conductors under strong magnetic fields: with the heat through the bulk of a quantum Hall bar being suppressed  [^3] Limitations of this picture have been discussed [286, 287]., the absence of backscattering leads to the heat to be transferred only between chirally connected reservoirs [142], as illustrated in Fig. 4(a).
@@
As this mechanism relies on the phase coherence of the electron waves, it is expected that the effect is reduced with the presence of dephasing mechanisms [150]. The addition of superconducting terminals confers a higher system tunability to the devices via the possibility to establish superconducting phase biases, but they do not seem to improve the performance of normal configurations.
Other operations related to heat rectification in which the current between two conductors is not reciprocal under swapping their temperatures have been proposed that rely on heat dissipated in the additional reservoirs [289, 290, 291, 292, 293, 148, 149, 294].
@@ V.3.2 Thermal transistors
Thermal gating of a particle current was observed in a system of capacitively coupled quantum dots [296], see Fig. 5(a) for an illustration. For low bias $|\Delta\mu|<U$, fluctuations in the occupation of the gate dot (which strongly depend on the temperature of its terminal) take the system dot in and out of the transport window, this way affecting the current by ±$\pm 30\,\mathrm{pA}$ [283]. The larger the temperature, the larger the modulation.
In order to avoid that the thermal gating involves a large transfer of heat, one needs to filter the conductor transitions [297, 293]. Consider the configuration described in Sec. IV.1, where the tunneling rates in the quantum dot $s$, $\Gamma_{n_{g}}^{i}$ ($i=L,R$), depend on the occupation, $n_{g}$, of quantum dot $g$. In the case that they are suppressed for a particular occupation of the dot (e.g., $\Gamma_{1}^{i}\ll\Gamma_{0}^{i}$), a fluctuation in the gate dot occupation switches on/off the transport through $s$. Furthermore, electrons entering $s$ will leave it with the same energy ($\varepsilon_{s}$, in this case): i.e., the current modulation does not require any heat transfer from the gate [297]. The necessary filtering can be done either with additional quantum dots [293] or using the spectral properties of the reservoirs. Using the gap of two superconducting terminals, the combination of two such thermal transistors at different temperatures was shown to enable a Maxwell demon [110]. These systems benefit from the energy exchange being quantized in unites of the electron-electron interaction, $U$.
An alternative is to consider the sensitivity of a quantum point contact to the charge of a nearby quantum dot [298] (the base): near pinch off, charging the base brings an open channel into the tunneling regime [299]. Bosonic base terminals have also been suggested [290, 260] which require more complicated setting to achieve the proper filtering.
@@ VI Discussion
The development of nanoscale devices with designed spectral properties gives access to the dynamics of few and even single particles and to manipulate them depending on their energy. This introduces mechanisms to control heat flows in quantum systems and use them for thermodynamic operations such as low power quantum heat engines, refrigerators or heat management devices. While their practical use in macroscopic setups is compromised by scalability and restricted to low temperatures, they could be incorporated to work onchip in quantum technologies, assisting on their energetics [300]. For this, multiterminal configurations able to separate the particle and heat flows are convenient versions of the thermocouple. We have reviewed different configurations of nanoscale energy harvesters, absorption and Peltier refrigerators and thermal circulators based in assorted configurations exploiting phase coherence, chirality, single-particle interactions an nonthermal states. Thermoelectric effects also allow to investigate properties that are opaque to conductance measurements such as the role of correlations or of mutual information.
@@ Acknowledgements.
We thank B. Sothmann for useful comments on the manuscript, and acknowledge collaborations over the years which led to some of the works discussed here, involving M. Acciai, B. Benítez, A. Braggio, P. Burset, M. Carrega, C. Elouard, G. Fleury, C. Gorini, D. Goury, G. Haack, F. Hajiloo, F. Haupt, G. Jaliel, A. N. Jordan, A. Levy Yeyati, R. López, G. Manzano, A. Mecha, L. W. Molenkamp, J. Monsel, P. P. Potts, R. K. Puddy, P. Samuelsson, J. Splettstoesser, D. Sánchez, C. G. Smith, B. Sothmann, S. M. Tabatabaei, L. Tesser, H. Thierschmann, R. S. Whitney. J. Yang, and the late M. Büttiker, with whom it was started. We acknowledge funding from the Spanish Ministerio de Ciencia e Innovación via grant No. PID2022-142911NB-I00, and through the “María de Maeztu” Programme for Units of Excellence in R&D CEX2023-001316-M.

### v2

@@ V Heat currents and cooling
The central part of this review has been devoted to nanoscale configurations able to generate electrical power in an isothermal conductor using some resource from a third one (usually heat). However, the same working principles have been proposed to operate these systems for a variety of different useful tasks. These include the nonlocal cooling of a cold reservoir, either using thermal (absorption refrigerators) or voltage (Peltier refrigerators) biases, and thermal management operations aimed at achieving control over heat flows [325]. We dedicate this last section to overview recent developments in nonlocal absorption and Peltier refrigerators and thermal management operation exclusive of multiterminal configurations: the circulation of heat and heat transistors.
@@ V.1 Absorption refrigerators
A multiterminal nanoscale system can work as an absorption refrigeration when there are at least two different temperatures among the electron reservoirs, so that under appropriate conditions the heat injected by the hottest terminal, $J_{H}>0$, forces the extraction of heat from the coldest one, $J_{C}>0$, in the absence of any electrochemical potential bias. In a three-terminal configuration with three temperatures, $T_{H}>T_{0}>T_{C}$, energy conservation ($J_{H}+J_{0}+J_{C}=0)$ and the second law of thermodynamics ($J_{H}/T_{H}-J_{0}/T_{0}-J_{C}/T_{C}\geq 0)$ impose that heat is dissipated in the reservoir with intermediate temperature, so we can identify $J_{H}$ as the only resource and define a coefficient of performance (COP) as: $\text{COP}=J_{C}/J_{H}$ , which is bounded from above by $\text{COP}_{C}=T_{C}/(T_{0}-T_{C})(1-T_{0}/T_{H})$.
@@
FIG. 6: Phase coherent absorption refrigerator. Heat is extracted from the cold reservoir of a three terminal system as the one depicted in Figure 3(a), with scattering regions 1 and 2 being quantum dots with resonance energies $\varepsilon_{1}$ and $\varepsilon_{2}$ and broadening $\Gamma=0.2k_{B}T$. (a) Cooling power, $J_{C}=J_{L}$, as a function of the resonance energies produced by a hot tip with $\Delta T_{H}/T=1$, and isothermal terminals $L$ and $R$, $T_{L}=T_{R}=T$. (b) Cooling power as a function of the heating at the tip, $\Delta T_{H}$, and of the right terminal, $\Delta T_{R}$, for $T_{L}=T$ and $(\varepsilon_{1}-\mu)/k_{B}T=2$, $(\varepsilon_{2}-\mu)/k_{B}T=4$. (c) Coefficient of performance for fixed $\Delta T_{H}=1$ but otherwise same parameters as in (b). Fixed parameters: $\mu=20k_{B}T$, tip transmission $\epsilon=0.5$, $d=2.5l_{0}$, $x=1.7l_{0}$, being $l_{0}=\hbar/\sqrt{8mk_{\text{B}}T}$, with the electron mass $m$.
@@
So far, Ref. [326] is the only absorption refrigerator in a noninteracting conductor to the best to our knowledge. The reason may be that this effect is very sensitive to the elastic coupling between terminals $L$ and $R$. We verified for this review that the same mechanism works for the fully phase-coherent implementation of the noninteracting resonant tunneling thermocouple discussed in Sec. III.2 and sketched in Fig. 3(a). The tip is used as a heat source that refrigerates one of the conductor terminals depending on its position. In Fig. 6(a) we show the heat flow $J_{L}$ when the tip temperature is increased by $\Delta T_{H}$ while maintaining $T_{L}=T_{R}=T$, as a function of the resonant level energies of the quantum dots. Cooling occurs for a very specific configuration where both levels are above the electrochemical potential, $\varepsilon_{\alpha}>\mu$, and destructive interference suppresses ${\cal T}_{LR}(E)$. In contrast to the inelastic version, no refrigeration is found in the case with $\varepsilon_{\alpha}<\mu$ (note that symmetries of the transmission probabilities exchanging $\varepsilon_{1}\leftrightarrow\varepsilon_{2}$ are only expected upon moving the tip from $x$ to dxd-x). The effect is also sensitive to an increase of the temperature of the other terminal, $\Delta T_{R}$, as shown in Fig. 6(b), as expected. The coefficient of performance is shown in Fig. 6(c) as a function of this temperature, attaining only a low fraction of the Carnot limit for this non optimized choice of parameters.
@@
A more robust effect is found due to interactions. Capacitively coupled quantum dots in the arrangement of Fig. 5(a) work as absorption refrigerators when the reservoir coupled to quantum dot $g$ is either the heat injector [96] or the refrigerated cold terminal [327, 328]. In the former case, it is sufficient that the heat conversion engine discussed in Sec. IV.1 operates in a configuration in which the generated current either extracts electrons from a reservoir ($L$ or $R$) above its electrochemical potential, or injects them below it [96]. Interestingly, the refrigeration of one conductor terminal can be combined with pumping heat on the hotter (but not hotter than the gate terminal) and additionally generating power, such that all currents in the conductor are reversed [193]. In the later case (cooling the terminal coupled to $g$), their operation relies in using the hot temperature of one of the terminals coupled to $s$ to favor the same series of fluctuations, $(n_{s},n_{g})=(0,0)\rightarrow(1,0)\rightarrow(1,1)\rightarrow(0,1)\rightarrow(0,0)$, cf. Eq. (21), that was used to transfer an energy quantum $U$ from $g$ to $s$ to power the heat engine. However, this requires a non trivial choice of the system parameters that limits its operating regime, and is restricted to low cooling powers.
@@
A photonic environment can also play the role of the hot resource for absorption refrigerators. Thermal photons coupled to a double quantum dot connecting two electronic reservoirs at different temperatures can drive processes that extract heat for the colder one [329, 330]. The inverse scenario, cooling a photonic resonator using two electronic reservoirs at different temperatures, was also explored in hybrid (superconductor-metal) configurations connected by a quantum dot [331].
@@ V.2 Nonlocal Peltier cooling
Analogously, the refrigeration side of chiral and Aharonov-Bohm engines discussed in Secs. III.3 and III.2, has been discussed using the split resonances of two quantum dots and taking advantage of the directionality imposed by the magnetic field [225, 337].
@@
In metallic islands, the energy balance needs to take the phonon bath into account, which adds an additional reservoir to the problem. Using the quasiparticle spectral properties of superconductor terminals in the current-carrying conductor, the temperature of a capacitively coupled single-electron box has been predicted to be reduced from $200\,\mathrm{mK}$ to $50\,\mathrm{mK}$ for islands of size $30\times 50\times 300\,\mathrm{nm^{3}}$ [340]. One can also extract heat from the phonon bath that couples to the nanostructure [173, 174], with similar ideas used for cooling of molecular vibrations close to their ground state [341, 342, 343, 344, 170, 345, 346, 347], as demonstrated in nanotube resonators [348].
@@ V.3.1 Heat circulators
The three terminal configuration also enables the possibility of defining circulation operations, where the heat injected from one terminal flows preferably in one direction (clock- or counterclockwise) to the next one, while transport in the opposite way is partially or totally blocked. A clear case where this happens is in chiral conductors under strong magnetic fields: with the heat through the bulk of a quantum Hall bar being suppressed  [^3] Limitations of this picture have been discussed [349, 350]., the absence of backscattering leads the heat to be transferred only between chirally connected reservoirs [76, 155], as illustrated in Fig. 4(a).
@@
As this mechanism relies on the phase coherence of the electron waves, it is expected that the effect is reduced with the presence of dephasing mechanisms [164]. The addition of superconducting terminals confers a higher system tunability to the devices via the possibility to establish superconducting phase biases [351], but they do not seem to improve the performance of normal configurations.
Other operations related to heat rectification in which the current between two conductors is not reciprocal under swapping their temperatures have been proposed that rely on heat dissipated in the additional reservoirs [352, 353, 354, 355, 356, 162, 163, 357, 358].
@@ V.3.2 Thermal transistors
Thermal gating of a particle current was observed in a system of capacitively coupled quantum dots [360], see Fig. 5(a) for an illustration. For low bias $|\Delta\mu|<U$, fluctuations in the occupation of the gate dot (which strongly depend on the temperature of its terminal) take the system dot in and out of the transport window, this way affecting the current by $\pm 30\,\mathrm{pA}$ [294]. The larger the temperature, the larger the modulation.
In order to avoid that the thermal gating involves a large transfer of heat, one needs to filter the conductor transitions [361, 356]. Consider the configuration described in Sec. IV.1, where the tunneling rates in the quantum dot $s$, $\Gamma_{n_{g}}^{i}$ ($i=L,R$), depend on the occupation, $n_{g}$, of quantum dot $g$. In the case that they are suppressed for a particular occupation of the dot (e.g., $\Gamma_{1}^{i}\ll\Gamma_{0}^{i}$), a fluctuation in the gate dot occupation switches on/off the transport through $s$. Furthermore, electrons entering $s$ will leave it with the same energy ($\varepsilon_{s}$, in this case): i.e., the current modulation does not require any heat transfer from the gate [361]. The necessary filtering can be done either with additional quantum dots [356] or using the spectral properties of the reservoirs. Using the gap of two superconducting terminals, the combination of two such thermal transistors at different temperatures was shown to enable a Maxwell demon [121]. These systems benefit from the energy exchange being quantized in units of the electron-electron interaction, $U$.
An alternative is to consider the sensitivity of a quantum point contact to the charge of a nearby quantum dot [362] (the base): near pinch off, charging the base brings an open channel into the tunneling regime [363]. Bosonic base terminals have also been suggested [353, 317, 330] which require more complicated settings to achieve the proper filtering.
@@ VI Discussion
The development of nanoscale devices with designed spectral properties gives access to the dynamics of few and even single particles and to manipulate them depending on their energy. This introduces mechanisms to control heat flows in quantum systems and use them for thermodynamic operations such as low power quantum heat engines, refrigerators or heat management devices. While their practical use in macroscopic setups is compromised by scalability and restricted to low temperatures, they could be incorporated to work onchip in quantum technologies, assisting on their energetics [364]. For this, multiterminal configurations able to separate the particle and heat flows are convenient versions of the thermocouple. We have reviewed different configurations of nanoscale energy harvesters, absorption and Peltier refrigerators and thermal circulators based in assorted configurations exploiting phase coherence, chirality, single-particle interactions an nonthermal states. Thermoelectric effects also allow to investigate properties that are opaque to conductance measurements such as the role of correlations or of mutual information, having hence and important role in fundamental research.
@@ Acknowledgements.
We thank J.-H. Jiang, J. Lu and B. Sothmann for useful comments on the manuscript, and acknowledge collaborations over the years which led to some of the works discussed here, involving M. Acciai, B. Benítez, A. Braggio, P. Burset, M. Carrega, C. Elouard, G. Fleury, C. Gorini, D. Goury, G. Haack, F. Hajiloo, F. Haupt, G. Jaliel, A. N. Jordan, A. Levy Yeyati, R. López, G. Manzano, A. Mecha, L. W. Molenkamp, J. Monsel, P. P. Potts, R. K. Puddy, P. Samuelsson, J. Splettstoesser, D. Sánchez, C. G. Smith, B. Sothmann, S. M. Tabatabaei, L. Tesser, H. Thierschmann, R. S. Whitney. J. Yang, and the late M. Büttiker, with whom it was started. We acknowledge funding from the Spanish Ministerio de Ciencia e Innovación via grants No. PID2022-142911NB-I00 and No. PID2024-157821NB-I0, and through the “María de Maeztu” Programme for Units of Excellence in R&D CEX2023-001316-M.